Using theorems, explain the causes of arbitrage between forward market hedge and money market hedge.
Answer #1

Interest rate parity:

As per Interest rate parity the difference in spot fate and forward rate exits due to differences in interest rate between two countries.

F/S = (1+ra)/(1+rb)
F= forward rate
S = spot rate.
ra = interest rate of price currency.
rb= interest rate of base currency.

If the above equation does not hold good then there is a chance of Arbitrage.

Purchasing power parity:

As per Purchasing power parity- the difference in spot fate and forward rate exits due to differences in inflation between two countries.


F/S = (1+Ia)/(1+Ib)
F= forward rate
S = spot rate.
Ia = inflation rate of price currency.
Ib= inflation rate of the base currency.

If the above equation does not hold good then there is a chance of Arbitrage.

the causes of arbitrage between forward market hedge and money market hedge is the deviation of should be interest rates or inflation rates from the actual interest rate or inflation rate.

In case of forward Market hedge: The person can exploit the opportunity by using the forward rate available in the market.

In case of money market hedge. The person uses the Interest rates and spot rate to exploit the opportunity.